A model rocket blasts off from the ground, rising straight upward with a constant acceleration that has a magnitude of 74.3 m/s2 for 1.80 seconds, at which point its fuel abruptly runs out. Air resistance has no effect on its flight. What maximum altitude (above the ground) will the rocket reach?


1
Expert's answer
2022-06-01T12:49:38-0400

Given:

"v_0=0, \; a=74.3\:{\rm m/s},\; t=1.80\:\rm s"


"h=h_1+h_2=v_0t+at^2/2+\frac{v_1^2}{2g}""v_1=v_0+at=at"

Hence, the maximum altitude that the rocket will reach

"h=at^2/2+\frac{(at)^2}{2g}"

"h=74.3*1.80^2/2+(74.3*1.80)^2/(2*9.81)\\
=1032\:\rm m"